A hammer of mass 1kg strikes on the head of the nail with the velocity of 2m/s.It drives the nail 0.01m into the wooden block.Find the force applied by the hammer and time of impact

Grade:11

1 Answers

Arun
6 years ago
Force applied by hammer= 200x1=200 Ntime of impact= 2/200=0.01 seconds=10 milli seconds2v=0 m/s , u=2 m/s => v^2-u^2 =2as => a=200 m/s